OpenStudy (anonymous):

The two figures are similar. The area of the smaller trapezoid is 771. The ratio of the smaller trapezoid to the bigger trapezoid is 28:64. What is the area of the bigger trapezoid? a) 4,096 b) 4,028 c) 784 d) 21

OpenStudy (anonymous):

Is there missing info from the question?

OpenStudy (anonymous):

Yes my bad! What is the area of the bigger trapezoid?

OpenStudy (anonymous):

Haha alright give me a min or two :)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

B :)

OpenStudy (anonymous):

Sorry, I took a bit longer than I thought. I was getting help on my own question.

OpenStudy (anonymous):

It's fine, thank you so much!

OpenStudy (anonymous):

No problem. Do you need to know how to do it, or are you okay?

OpenStudy (anonymous):

Can you please explain it. :)

OpenStudy (anonymous):

Sure thing!

OpenStudy (kropot72):

@kristinalgarcia I already showed you how to do basically the same question.

OpenStudy (anonymous):

Yes I'm aware of that but when I did it with the way you showed me on this problem the answer I got was not an answer choice.

OpenStudy (anonymous):

Okay so they gave you the are of the *similar* trapezoid which was 771. And the ratio was 28:64. So, 28/28=1 , and 64/28 = 2.28571429... When you square 2.28... : (2.285...)^2 , you get 5.2244... which is the ratio of the areas of the similar trapezoids. So, the area of the bigger trapezoid is 5.2244... times bigger than the smaller one: 771(5.2244...) = 4,028.0816... => 4,028

OpenStudy (kropot72):

Just multiply by the square of the scale factor: \[\large 771\times(\frac{64}{28})^{2}=you\ can\ calculate\]

OpenStudy (anonymous):

Now I see what I did wrong. @kropot72

OpenStudy (kropot72):

@kristinalgarcia The way that I showed you gives one of the answer choices.

OpenStudy (anonymous):

Yep, ^ that is an easier, much more efficient way of solving it, provided you have a graphing calculator or sufficient calculator app. I can't do that on my phone, so I did it an alternate way, but both will give you the same answer! :)

OpenStudy (anonymous):

@Thatsodan Thank you!

OpenStudy (anonymous):

Of course!